An automatic machine fills jars of jam with a mean net weight of 340 grams. Assume a normal distribution with a standard deviation of 8 grams. What is the probability that a randomly selected jar of jam weighs between 338 grams and 344 grams
1
Expert's answer
2019-04-18T10:43:20-0400

P(338<X<344)=P((338-340)/8<Z<(344-340)/8)=P(-0.25<Z<0.5)=

=P(Z<0.5)-P(Z<-0.25)=P(Z<0.5)-1+P(Z<0.25)=0.6915-1+0.5987=0.2902.